# Does rare, noise-induced, bypass transition in plane Couette flow bypass instantons ?

### v1

@@ 4.1 Probability of crossing and mean waiting time
We present the probability to observe a build up computed by means of AMS and DNS in figure 13 (a). The corresponding mean first passage time before crossing to turbulence, computed solely by means of AMS, is displayed in figure 13 (b). Both are presented in logarithmic scale. These two quantities indicate that the noise induced transition becomes rapidly very rare as $\beta$ is increased: the estimate of $\log(\alpha)$ decreases by more than five decades, while that of $\log(T)$ increases by five decades in the range of $\beta$ considered here. This corresponds to variations by more than twenty orders of magnitude in linear scale. Most of this rapid variation takes place between $\beta=2.7\cdot 10^{5}$ and $\beta=6\cdot 10^{5}$. In that range both $\log(\alpha)$ and $\log(T)$ are very well approximated by affine functions. The fit are added to figure 13 (a,b) and respectively give the slopes $(-1.53\pm 0.01)\cdot 10^{-4}$ and $(1.53\pm 0.01)\cdot 10^{-4}$. This confirms that the probability $\alpha$ decays exponentially and that the mean first passage time $T$ increases exponentially and that their rate are opposite of one another. This scaling of $T$ and $\alpha$ in $\beta$ is the same as the one observed in most systems where rare bistability is present.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 13: From computations in system of size $L_{x}\times L_{z}=6\times 4$ at Reynolds number $R=500$ under divergence free forcing, by means of AMS and up to $\beta=3\cdot 10^{5}$ by means of DNS. (a) : Probability to cross from the defined neighbourhood of the laminar baseflow to the turbulent flow as a function of $\beta$, an affine fit is performed for $\beta\geq 2.7\cdot 10^{5}$ and is added to the plot. (b) : Mean first passage time before transition to the turbulent flow as a function of $\beta$, an affine fit is performed for $\beta\geq 2.7\cdot 10^{5}$ and is added to the plot. (c) : average duration of reactive trajectories as a function of $\beta$. (d) : normalised standard deviation of the estimate of $\alpha$ as a function of $\beta$, the black dashed line indicates the ideal value $1$.
We perform two tests to validate the computation of $\alpha$. Up to $\beta=3\cdot 10^{5}$, computations by means of DNS are conceivable since $\alpha\geq 10^{-3}$. The results are added to figure 13 (a) and show an overlap between the AMS and DNS intervals of confidence. This indicates that the values computed by means of AMS can be trusted in that range. This indicates that the exponential decrease of $\alpha$ is valid in that range. We complement this test by the computation of $\sigma_{\rm norm}$, the normalised variance of the estimate of $\alpha$ (see Eq. (35), § C), which is presented in figure 13 (d). In the whole range of interest, we can see that this quantity is around one for the whole range of $\beta$ considered here. This confirms that the estimates performed by means of AMS are not tainted by an uncontrolled error, and that the exponential decrease of $\alpha$ is valid in the whole range of interest. Further tests of validity of computations using this indicator are presented in appendix C.
@@ 4.2 Duration of trajectories
Refer to captionRefer to caption
Figure 14: (a) Logarithm of the duration of reactive trajectories as a function of $\beta$ for various forcing shapes (spectrum, resolution, forced components) computed by means of AMS in the MFU size system. (b) Logarithm of these duration of reactive trajectories as a function of $1/\epsilon$ for various forcing shapes (spectrum, resolution, forced components).
@@ 4.2.1 Forcing on components
The average duration $\tau$ of build up trajectories is computed as a function of $\beta$ for several shape of the spectrum of the forcing on components ($\sigma_{y,z}=0$ or not, several values of $N_{r}$ at several resolutions). We first observe the logarithm of $\tau$ as a function of $\beta$, for all these cases in figure 14, (a). We note that, as is commonly reported, $\tau$ increases with $\beta$ at fixed spectrum shape. Moreover, the growth of the logarithm of $\tau$ is faster than linear with the parameter $\beta$ controlling the energy injection rate independently of the structure of the forcing. We can observe this plot differently and consider $\tau$ at fixed $\beta$ as the spectrum shape is changed. In particular, we consider the effect of a decreased energy injection rate as less modes are forced (as indicated in Eq. 8). Among forcing on all three components, we note that as we decrease $N_{r}$ from $8$ to $4$, thus reducing the range of low wavenumbers where the spectrum is white and consequently reducing $\epsilon$, $\tau$ is larger and larger. At fixed $N_{r}=4$, as we go from forcing on three components to forcing on the one streamwise component, $\tau$ is even larger. From these two observations, we can conclude that our estimates of $\tau$ all increase as the energy injection rate $\epsilon$ is decreased: indeed, equation (8) indicates that both increasing $\beta$ and decreasing $N_{r}$ reduces the energy injection rate. We now test this dependence even further and ask whether $\tau$ is solely a function of the energy injection rate in this forcing family. For this matter, we display $\tau$ as a function of $1/\epsilon$ (Fig. 14, (b)). The curves of $\log(\tau)\left(\frac{1}{\epsilon}\right)$ are regrouped when we use this rescaling. Moreover, four of the six cases tested here, which have the same numerical resolution, collapse together. As such, we have a behaviour similar to what had been observed on the averaged kinetic energy conditioned on no transition to occur (§ 2.2.2, Fig. 4). The matter is not trivial: we expect $\tau$ to generally grow with $\frac{1}{\epsilon}$: more energy injected per unit time should mean faster travel toward turbulence. However, there is also no reason why the shape of the spectrum of the forcing noise should have no subtle effect, we have not examined the energy transfer, direct and indirect, in detail. From these AMS computations, we also estimated the probability of crossing and the mean first passage time before crossing (not shown here). Similarly to what has been observed in the former section, we have respectively a rapid decrease of $\alpha$ and a rapid increase of $T$ as $\beta$ is increased. Moreover, we also note that the effect of the spectrum shape is similar to what has been observed with $\tau$ : at fixed $\beta$, injecting less energy leads to a decrease of $\alpha$ and an increase of $T$. Observing those quantities as a function of $1/\epsilon$ regroups the plots together, but as tightly as for $\tau$.
@@
    Firstly, since we find the same dependence of average kinetic energy, the same type of path (in term of stages) and the same type of dependence of their duration on $\epsilon$, regardless of the forcing structure within this forcing family, we can further back the assertion that the computed reactive trajectories are selected by the non linear dynamics and not by the forcing. Behind these tests, we aim to investigate the question of universality in these paths: to what extent are they a product of the Navier–Stokes equations alone and the energy injection rate as a sole parameter, and to what extent do they depend on the forcing spectrum.
@@
    Secondly, the growth of $\tau$ with the parameter controlling the noise variance (or energy injection rate) is apparently exponential: its logarithm is convex. As such it is quite faster then what had been previously observed in simpler systems [Rolland and Simonnet, 2015, Rolland et al., 2016]. Indeed, it has been demonstrated in one degree of freedom systems having a simple saddle between both metastable states [Cérou et al., 2013] and verified numerically in such systems having more DoF [Rolland and Simonnet, 2015, Rolland et al., 2016] that the duration of reactive trajectories grew logarithmically with the variance of the noise forcing. When the saddle between both states become extremely flat, the reactive trajectories undergoes a random walk on the plateau between both states. In that case, a faster linear increase of the duration of reactive trajectories is measured [Rolland et al., 2016]. However, this leads to a Gaussian distribution of trajectory durations, unlike the Gumbel observed here (Fig. 12). An exponential increase can be seen in simple models if there is a local (weakly) metastable minimum standing between the two main metastable states. In that case, the exponential increase is caused by the waiting time before escape of the secondary metastable state [Rolland and Simonnet, 2015]. In the case of build up of turbulence, there is no known intermediate metastable state. However the complexity of phase space could play a similar role.
@@ 4.2.2 Divergence free forcing
We present the average duration of reactive trajectories obtained under divergence free forcing as a function of $\beta$ in figure 13 (c). In contrast with the behaviour of reactive trajectories obtained with forcing on components, the average duration of reactive trajectories grows rather slowly. Indeed, we can observe an increase of $\tau$ with $\beta$ up to $\beta\lesssim 260000$, followed by a plateau for $260000\lesssim\beta\lesssim 3300000$. Above that value of $\beta\gtrsim 330000$, the average duration of reactive trajectories estimated by means of AMS is smaller. As such, it is not straightforward to precisely fit a known functional dependence of $\tau(\beta)$ on that data, even if the slow logarithmic growth would be the closest to what we estimated. There are apparent distinctions with that behaviour, such as the local maximum. Such maxima are not commonly reported in simpler noise driven transitions. An effect of the large error bars cannot be ruled out, especially since no change of tendency is observed in $\log(\alpha)$, $\log(T)$ nor in the trajectories.
@@ 5.2 Mechanical description of the paths
Observation of the trajectories indicated that no matter the domain size and the structure of forcing, the flow responded to the added noise by generating streamwise invariant streamwise velocity tubes, along with weak streamwise vorticity. On the trajectory, these tubes gradually accumulate energy obtained from the forcing, in that way the non trivial flow is sustained by the statistically steady energy injection. The tubes finally become intense enough for the self sustaining process of turbulence to start, after the trajectory has crossed the separatrix between laminar and turbulent flow. It is notable that plane Couette flow was efficiently able to generate these tubes, even if the wavenumber forcing range was distinct from the natural wavenumber range of transitional turbulence. In that respect, these trajectories are distinct from the one where the wavenumber of velocity streaks and streamwise vortices are directly non-linearly forced [Rigas et al., 2021]. This indicates that the flow is able to redistribute the energy from one range of wavenumbers to another, even in the regime where the energy injection rate is small. Moreover, this hints that the resulting coherent structures (in the form of streamwise velocity tubes) are naturally selected by plane Couette flow, not imposed by the forcing. In that respect, this would mean that we observed a non linear receptivity of the flow in the first part of the trajectory. Because that receptivity was independent from the several types of forcing we applied, it could indicate some universality in the response of the flow to the forcing. That search for universal features in the transition, such as a mediator flow state arising solely from the Navier–Stokes equations in plane Couette flow, and the corresponding mechanical processes, is then in line with the computation of the properties of the separatrix such as the saddle on it or the point on it which is the closest to the laminar flow, as measured by kinetic energy. The build up is distinct from these last two flow states because minimal seeds are precisely defined non linear flows and would not naturally emerge from finite variance noise and, as we will further discuss, edge states are avoided by the computed build up trajectories. What we observe here is closer to (non)-linear response, with the distinction that we have not optimised the forcing, and we are not observing the optimal response to any regular enough forcing.
@@
Similarly to paths to turbulence identified using other approaches (forcing, bisections, optimisations), the reactive trajectories computed here display a spatial localisation of the triggered velocity streaks as the domain size is increased. This localisation is visible in the last state at the last stage. The main difference is that in those flow states the actual self sustaining process of turbulence is triggered in a narrow region of the domain.
@@ 5.3 Statistical description of the paths
On several aspects, there is a similarity between the trajectories of build up of turbulence under forcing and that of collapse of turbulence. The paths followed in the plane $\log(E_{k,x})-\log(E_{k,y-z})$ both show a strong convexity. The succession of events : first the generation of streamwise velocity tubes followed by the appearance of streamwise vorticity (with spanwise localisation when domain size is increased) can qualitatively seem as a rewind of those observed in the collapse of turbulence [Rolland, 2022, Liu et al., 2021, Gomé et al., 2022]. Elements of this scenario had been noted in earlier studies without being at the centre of attention yet [Philip and Manneville, 2011, Chantry and Schneider, 2014]. However, the build-up and collapse trajectories are by no mean time reversed paths of one another. Firstly the aforementioned similarity is not exact. Secondly, they take place under clearly distinct conditions: the build up is forced, while collapses are purely deterministic. A corollary is that build up paths cross a separatrix, with the help of the forcing, while the collapse paths do not: they avoid it. This induces further differences, such as the duration of each phase: the generation of the streamwise vorticity tubes shows a polynomial time dependence, while the decay of said tube is exponential in time [Rolland, 2022]. The distinction between the forward and backward types of paths is also different from what is found in noise driven non-gradient system. In those system, there is a clearly distinction between the backward and forward path. However, they both comprise of a fluctuation path toward the same saddle, followed by a relaxation path. Meanwhile, the build up has a fluctuation path toward the separatrix then relaxes, while the collapse can be viewed as a pure relaxation path.
@@
The comparison between the reactive trajectory and the nearest fixed point on the crossed separatrix is very often the focus of attention. This is because, as stated earlier, in systems which do not have a degenerate saddle [Bouchet and Touchette, 2012], in the limit of energy injection and/or forcing noise variance going to zero, the reactive trajectories are structured by instantons, which cross the separatrix right at a saddle point (see the sketch in figure 15 (a)). In said regime, one finds a structuration of the bistability: an exponential dependence of the mean first passage time and crossing probability in the noise variance, a concentration of trajectories in phase space and a specific distribution of reactive trajectories duration, to mention the most important features. Surprisingly enough, all these properties were found in our reactive trajectories.
@@
    The paths sampled by means of AMS could be incorrect. This can happen when the reaction coordinate is not adapted. One then observes the apparent bias phenomenon: most sample reactive trajectories do not follow the most likely path [Bréhier et al., 2016, Lucente et al., 2022]. However, that type of behaviour would show a clear distinction between the trajectory durations computed by means of AMS and those computed by means of DNS as well as lead to a very large relative sample variance of the probability of crossing, which is not the case here. Moreover, when this phenomenon manifests itself, the reactive trajectories tend to select a path that still visit the neighbourhood of a saddle [Rolland and Simonnet, 2015].
@@
    One can observe correctly computed reactive trajectories which are outside of the instanton regime if the relevant saddle is very flat [Rolland et al., 2016]. In that situation, one does observes an exponential growth of mean first passage time and an exponential decay of the crossing probability, which is our case. Meanwhile, one observes a Gaussian distribution of trajectory duration and a linear increase of their average duration and variance, which is not our case. Moreover, this regime still visits the neighbourhood of a saddle, which is not our case.
@@
    When the saddle in the system can become degenerate, one can observe sub-instantons, that do not follow the expected path [Bouchet and Touchette, 2012]. However, in that case, the probability of crossing and the mean first passage time have a very distinct scaling in exponential of the square root of the noise variance, which is not our case.

### v2

@@ 4.1 Probability of crossing and mean waiting time
Refer to caption Refer to caption Refer to caption
Figure 14: From computations in system of size $L_{x}\times L_{z}=6\times 4$ at Reynolds number $R=500$ under divergence free forcing, by means of AMS and up to $\beta=3\cdot 10^{5}$ by means of DNS. (a) : Probability to cross from the defined neighbourhood of the laminar baseflow to the turbulent flow as a function of $\beta$, an affine fit is performed for $\beta\geq 2.7\cdot 10^{5}$ and is added to the plot. (b) : Mean first passage time before transition to the turbulent flow as a function of $\beta$, an affine fit is performed for $\beta\geq 2.7\cdot 10^{5}$ and is added to the plot. (c) : normalised standard deviation of the estimate of $\alpha$ as a function of $\beta$, the black dashed line indicates the ideal value $1$.
Figure 15: From the same AMS computations as in figure 14 average duration of reactive trajectories as a function of $\beta$.
@@
We present the probability to observe a build up computed by means of AMS and DNS in figure 14 (a). The corresponding mean first passage time before crossing to turbulence, computed solely by means of AMS, is displayed in figure 14 (b). Both are presented in logarithmic scale. These two quantities indicate that the noise induced transition becomes rapidly very rare as $\beta$ is increased: the estimate of $\log(\alpha)$ decreases by more than five decades, while that of $\log(T)$ increases by five decades in the range of $\beta$ considered here. This corresponds to variations by more than twenty orders of magnitude in linear scale. Most of this rapid variation takes place between $\beta=2.7\cdot 10^{5}$ and $\beta=6\cdot 10^{5}$. In that range both $\log(\alpha)$ and $\log(T)$ are very well approximated by affine functions. The fit are added to figure 14 (a,b) and respectively give the slopes $(-1.53\pm 0.01)\cdot 10^{-4}$ and $(1.53\pm 0.01)\cdot 10^{-4}$. This confirms that the probability $\alpha$ decays exponentially and that the mean first passage time $T$ increases exponentially and that their rate are opposite of one another. This scaling of $T$ and $\alpha$ in $\beta$ is the same as the one observed in most systems where rare bistability is present. In that framework, as we had stated in section 1.1, the decay rate of the probability $\alpha$ and the growth rate of the mean first passage time $T$ is formally given by the difference of quasipotential between the visited saddle and the first multistable state Bouchet and Reygner, (2016). In our case, this would correspond to the edge state and the laminar baseflow. Note however that the computation of the quasipotential and the estimation of its value at specific flow states is beyond the scope of this work.
We perform two tests to validate the computation of $\alpha$. Up to $\beta=3\cdot 10^{5}$, computations by means of DNS are conceivable since $\alpha\geq 10^{-3}$. The results are added to figure 14 (a) and show an overlap between the AMS and DNS intervals of confidence. This indicates that the values computed by means of AMS can be trusted in that range. This indicates that the exponential decrease of $\alpha$ is valid in that range. We complement this test by the computation of $\sigma_{\rm norm}$, the normalised variance of the estimate of $\alpha$ (see Eq. (23), § D), which is presented in figure 14 (c). In the whole range of interest, we can see that this quantity is around one for the whole range of $\beta$ considered here. This confirms that the estimates performed by means of AMS are not tainted by an uncontrolled error, and that the exponential decrease of $\alpha$ is valid in the whole range of interest. Further tests of validity of computations using this indicator are presented in appendix D.
@@ 4.2 Duration of trajectories
Figure 16: Logarithm of the duration of reactive trajectories as a function of $\beta$ for various forcing shapes (spectrum, resolution, forced components) computed by means of AMS in the MFU size system.
@@ 4.2.1 Forcing on components
The average duration $\tau$ of build up trajectories is computed as a function of $\beta$ for several shape of the spectrum of the forcing on components ($\sigma_{y,z}=0$ or not, several values of $N_{r}$ at several resolutions). We first observe the logarithm of $\tau$ as a function of $\beta$, for all these cases in figure 16, (a). We note that, as is commonly reported, $\tau$ increases with $\beta$ at fixed spectrum shape. Moreover, the growth of the logarithm of $\tau$ is faster than linear with the parameter $\beta$ controlling the energy injection rate independently of the structure of the forcing. We can observe this plot differently and consider $\tau$ at fixed $\beta$ as the spectrum shape is changed. In particular, we consider the effect of a decreased energy injection rate as less modes are forced (as indicated in Eq. 8). Among forcing on all three components, we note that as we decrease $N_{r}$ from $8$ to $4$, thus reducing the range of low wavenumbers where the spectrum is white and consequently reducing $\epsilon$, $\tau$ is larger and larger. At fixed $N_{r}=4$, as we go from forcing on three components to forcing on the one streamwise component, $\tau$ is even larger. From these two observations, we can conclude that our estimates of $\tau$ all increase as the energy injection rate $\epsilon$ is decreased: indeed, equation (8) indicates that both increasing $\beta$ and decreasing $N_{r}$ reduces the energy injection rate. From these AMS computations, we also estimated the probability of crossing and the mean first passage time before crossing (not shown here). Similarly to what has been observed in the former section, we have respectively a rapid decrease of $\alpha$ and a rapid increase of $T$ as $\beta$ is increased. Moreover, we also note that the effect of the spectrum shape is similar to what has been observed with $\tau$ : at fixed $\beta$, injecting less energy leads to a decrease of $\alpha$ and an increase of $T$. We have observed as a function of $\epsilon$ the average duration of trajectories $\tau$, the probability of observing a trajectory $\alpha$ and the mean first passage time before a trajectory occurs (not shown here). While this rescaling tend to regroup the curves, the precise collapse is debatable. The matter is not trivial: we expect $\tau$ to generally grow with $\frac{1}{\epsilon}$: more energy injected per unit time should mean faster travel toward turbulence. However, there is also no reason why the shape of the spectrum of the forcing noise should have no subtle effect, we have not examined the energy transfer, direct and indirect, in detail.
@@
    Firstly, we find that the energy injection rate $\epsilon$ has the same type of effect on the average kinetic energy, on the type of path generated (in term of stages) and on quantitative properties of the paths like their average duration $\tau$, with moderate effect of the forcing structure within this forcing family. We can further back the assertion that the computed reactive trajectories are selected by the non linear dynamics and not by the forcing spectrum. Behind these tests, we aim to investigate the question of universality in these paths: to what extent are they a product of the Navier–Stokes equations alone and the energy injection rate as a sole parameter, and to what extent do they depend on the forcing spectrum.
@@
    Secondly, the growth of $\tau$ with the parameter controlling the noise variance (or energy injection rate) is apparently exponential: its logarithm is convex. As such it is quite faster then what had been previously observed in simpler systems Rolland and Simonnet, (2015); Rolland et al., (2016). Indeed, it has been demonstrated in one degree of freedom systems having a simple saddle between both metastable states Cérou et al., (2013) and verified numerically in such systems having more DoF Rolland and Simonnet, (2015); Rolland et al., (2016) that the duration of reactive trajectories grew logarithmically with the variance of the noise forcing. When the saddle between both states become extremely flat, the reactive trajectories undergoes a random walk on the plateau between both states. In that case, a faster linear increase of the duration of reactive trajectories is measured Rolland et al., (2016). However, this leads to a Gaussian distribution of trajectory durations, unlike the Gumbel observed here (Fig. 13). An exponential increase can be seen in simple models if there is a local (weakly) metastable minimum standing between the two main metastable states. In that case, the exponential increase is caused by the waiting time before escape of the secondary metastable state Rolland and Simonnet, (2015). In the case of build up of turbulence, there is no known intermediate metastable state. However the complexity of phase space could play a similar role.
@@ 4.2.2 Divergence free forcing
We present the average duration of reactive trajectories obtained under divergence free forcing as a function of $\beta$ in figure 15. In contrast with the behaviour of reactive trajectories obtained with forcing on components, the average duration of reactive trajectories grows rather slowly. Indeed, we can observe an increase of $\tau$ with $\beta$ up to $\beta\lesssim 260000$, followed by a plateau for $260000\lesssim\beta\lesssim 3300000$. Above that value of $\beta\gtrsim 330000$, the average duration of reactive trajectories estimated by means of AMS is smaller. As such, it is not straightforward to precisely fit a known functional dependence of $\tau(\beta)$ on that data, even if the slow logarithmic growth would be the closest to what we estimated. There are apparent distinctions with that behaviour, such as the local maximum. Such maxima are not commonly reported in simpler noise driven transitions. An effect of the large error bars cannot be ruled out, especially since no change of tendency is observed in $\log(\alpha)$, $\log(T)$ nor in the trajectories.
@@ 5.2 Mechanical description of the paths
Observation of the trajectories indicated that no matter the domain size and the structure of forcing, the flow responded to the added noise by generating streamwise invariant streamwise velocity tubes, along with weak streamwise vorticity. On the trajectory, these tubes gradually accumulate energy obtained from the forcing, in that way the non trivial flow is sustained by the statistically steady energy injection. The tubes finally become intense enough for the self sustaining process of turbulence to start, after the trajectory has crossed the separatrix between laminar and turbulent flow. It is notable that plane Couette flow was efficiently able to generate these tubes, even if the wavenumber forcing range was distinct from the natural wavenumber range of transitional turbulence. In that respect, these trajectories are distinct from the one where the wavenumber of velocity streaks and streamwise vortices are directly non-linearly forced Rigas et al., (2021). This indicates that the flow is able to redistribute the energy from one range of wavenumbers to another, even in the regime where the energy injection rate is small. Moreover, this hints that the resulting coherent structures (in the form of streamwise velocity tubes) are naturally selected by plane Couette flow, not imposed by the forcing. In that respect, this would mean that we observed a non linear receptivity of the flow in the first part of the trajectory. Because that receptivity was independent from the several types of forcing we applied, it could indicate some universality in the response of the flow to the forcing. That search for universal features in the transition, such as a mediator flow state arising solely from the Navier–Stokes equations in plane Couette flow, and the corresponding mechanical processes, is then in line with the computation of the properties of the separatrix such as the saddle on it or the point on it which is the closest to the laminar flow, as measured by kinetic energy. The build up is distinct from these last two flow states because minimal seeds are precisely defined non linear flows and would not naturally emerge from finite variance noise and, as we will further discuss, edge states are avoided by the computed build up trajectories. What we observe here is closer to (non)-linear response, with the distinction that we have not optimised the forcing, and we are not observing the optimal response to any regular enough forcing.
@@
Similarly to paths to turbulence identified using other approaches (forcing, bisections, optimisations), the reactive trajectories computed here display a spatial localisation of the triggered velocity streaks as the domain size is increased. This localisation is visible in the last state at the last stage. The main difference is that in those flow states the actual self sustaining process of turbulence is triggered in a narrow region of the domain. We have been able to observe spanwise localisation in domains of size $L_{x}\times L_{z}=18\times 12$ and $L_{x}\times L_{z}=36\times 24$. We expect that with further increased domain size, such as $L_{x}\times L_{z}=68\times 34$, we will also observe streamwise localisation of triggered turbulence, as it is the case in edge states and minimal seeds. We have not performed these computation for this text due to the increased numerical cost. As explained in appendix D, we expect that AMS computations in such a larger domain would function as well as those presented in this text.
@@ 5.3 Statistical description of the paths
On several aspects, there is a similarity between the trajectories of build up of turbulence under forcing and that of collapse of turbulence. The paths followed in the plane $\log(E_{k,x})-\log(E_{k,y-z})$ both show a strong convexity. The succession of events : first the generation of streamwise velocity tubes followed by the appearance of streamwise vorticity (with spanwise localisation when domain size is increased) can qualitatively seem as a rewind of those observed in the collapse of turbulence Rolland, (2022); Liu et al., (2021); Gomé et al., (2022). Elements of this scenario had been noted in earlier studies without being at the centre of attention yet Philip and Manneville, (2011); Chantry and Schneider, (2014). However, the build-up and collapse trajectories are by no mean time reversed paths of one another. Firstly the aforementioned similarity is not exact. Secondly, they take place under clearly distinct conditions: the build up is forced, while collapses are purely deterministic. A corollary is that build up paths cross a separatrix, with the help of the forcing, while the collapse paths do not: they avoid it. This induces further differences, such as the duration of each phase: the generation of the streamwise vorticity tubes shows a polynomial time dependence, while the decay of said tube is exponential in time Rolland, (2022). The distinction between the forward and backward types of paths is also different from what is found in noise driven non-gradient system. In those system, there is a clearly distinction between the backward and forward path. However, they both comprise of a fluctuation path toward the same saddle, followed by a relaxation path. Meanwhile, the build up has a fluctuation path toward the separatrix then relaxes, while the collapse can be viewed as a pure relaxation path.
@@
The comparison between the reactive trajectory and the nearest fixed point on the crossed separatrix is very often the focus of attention. This is because, as stated earlier, in systems which do not have a degenerate saddle Bouchet and Touchette, (2012), in the limit of energy injection and/or forcing noise variance going to zero, the reactive trajectories are structured by instantons, which cross the separatrix right at a saddle point (see the sketch in figure 17 (a)). In said regime, one finds a structuration of the bistability: an exponential dependence of the mean first passage time and crossing probability in the noise variance, a concentration of trajectories in phase space and a specific distribution of reactive trajectories duration, to mention the most important features. Surprisingly enough, all these properties were found in our reactive trajectories.
@@
    The paths sampled by means of AMS could be incorrect. This can happen when the reaction coordinate is not adapted. One then observes the apparent bias phenomenon: most sample reactive trajectories do not follow the most likely path Bréhier et al., (2016); Lucente et al., (2022). However, that type of behaviour would show a clear distinction between the trajectory durations computed by means of AMS and those computed by means of DNS as well as lead to a very large relative sample variance of the probability of crossing, which is not the case here. Moreover, when this phenomenon manifests itself, the reactive trajectories tend to select a path that still visit the neighbourhood of a saddle Rolland and Simonnet, (2015).
@@
    One can observe correctly computed reactive trajectories which are outside of the instanton regime if the relevant saddle is very flat Rolland et al., (2016). In that situation, one does observes an exponential growth of mean first passage time and an exponential decay of the crossing probability, which is our case. Meanwhile, one observes a Gaussian distribution of trajectory duration and a linear increase of their average duration and variance, which is not our case. Moreover, this regime still visits the neighbourhood of a saddle, which is not our case.
@@
    When the saddle in the system can become degenerate, one can observe sub-instantons, that do not follow the expected path Bouchet and Touchette, (2012). However, in that case, the probability of crossing and the mean first passage time have a very distinct scaling in exponential of the square root of the noise variance, which is not our case.